OpenStudy (anonymous):

Solve for x: ln(3x)=12

OpenStudy (anonymous):

Divide both sides by three you you got your answer. : )

OpenStudy (anonymous):

ln(x)=4? then e^ln(x)=e^4 -> x=e^4 ?

OpenStudy (anonymous):

the number 4 goes into the number 12. So 3*(4)=12

OpenStudy (anonymous):

or is it e^ln(3x)=e^12 -> 3x=e^12-> x=e^12/3 ?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

\[\ln(3x)=12\iff e^{12}=3x\]

OpenStudy (anonymous):

and so \[x=\frac{e^{12}}{3}\]

OpenStudy (anonymous):

Okay, @satellite73. Thanks! :D

OpenStudy (anonymous):

lol yw you did all the work

OpenStudy (anonymous):

haha, still... Had to know if i was right :P